# Underwater Variable Zoom: Depth-Guided Perception Network for Underwater Image Enhancement

### v1

@@ Abstract
Underwater scenes intrinsically involve degradation problems owing to heterogeneous ocean elements. Prevailing underwater image enhancement (UIE) methods stick to straightforward feature modeling to learn the mapping function, which leads to limited vision gain as it lacks more explicit physical cues (e.g., depth). In this work, we investigate injecting the depth prior into the deep UIE model for more precise scene enhancement capability. To this end, we present a novel depth-guided perception UIE framework, dubbed underwater variable zoom (UVZ). Specifically, UVZ resorts to a two-stage pipeline. First, a depth estimation network is designed to generate critical depth maps, combined with an auxiliary supervision network introduced to suppress estimation differences during training. Second, UVZ parses near-far scenarios by harnessing the predicted depth maps, enabling local and non-local perceiving in different regions. Extensive experiments on five benchmark datasets demonstrate that UVZ achieves superior visual gain and delivers promising quantitative metrics. Besides, UVZ is confirmed to exhibit good generalization in some visual tasks, especially in unusual lighting conditions. The code, models and results are available at: https://github.com/WindySprint/UVZ.
Index Terms:
Underwater image enhancement, Depth-guided perception, Near-far scenarios, Color correction.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
In the early days of UIE field, researchers have proposed model-free and model-based traditional methods. Model-free methods [5, 15, 16] enhances underwater images by modifying pixels in degraded channels and blurred regions. On the other hand, model-based methods [7, 17] introduce depth information and models to build the underwater degradation process, and reverse it to restore underwater images. In recent years, deep learning methods [11, 18, 19] have shown remarkable enhancements, leveraging benchmark data and miscellaneous networks. Despite significant progress achieved with the above methods, challenges emerge when addressing widely varying degrees of regional degradation. Fig. 1 exhibits the enhancement effects of some UIE methods on near-far scenarios. It is evident that the degradation degree varies between the near scene (orange region) and far scene (red region) in the raw image. In comparison, the visual quality of the former and the adjacent region still exhibits relatively better performance. While most methods have considerable visual improvement in the near scene, but still retain notable color deviation in the far scene. For more accurate color enhancement, we advocate that this process should incorporate depth information related to scene distance, which allows for a more precise division of different regions in underwater images.
Previous works [12, 13, 20] have incorporated depth maps to explore physical cues, contributing to the response of degraded regions. However, direct weight adjustments may restrict the exploration of visual cues, as the adjacent features of far and near points exhibit varying gains in the enhancement process. With the above analysis, we hope to construct an adaptive region enhancement method. Before enhancing each point, the depth information should be used to determine the importance of non-local and local information. In the near-field region, local information is critical to the quality of the enhanced image, whereas in the far-field region, non-local information can provide valuable insights.
@@
To achieve adaptive learning in the near-far region, this paper proposes a novel UIE method based on depth-guided perception, UVZ, which includes depth estimation network (DEN), auxiliary supervision network (ASN), and depth-guided enhancement network (DGEN). First, the proposed model generates a depth map by DEN, and uses ASN to further supervise the rationality of the depth map during training. Meanwhile, the desired features for different regions are captured by non-local and local branches in DGEN. To adapt to varying receptive fields, we perform a R3N transformation on the estimated depth map, guiding feature fusion in different regions. Through the above process, UVZ effectively and comprehensively enhances the color, contrast, and clarity of underwater images.
@@
• We designed three sub-networks to compose a new depth-guided perception framework, UVZ. For non-homogeneous degraded underwater scenes, we employs a learning strategy that adaptively combines non-local and local features.
@@
• To enhance the ability to distinguish degraded regions, we incorporate an attention estimation-scene reconstruction training process. This step aims to strengthen the correlation between the depth map and the underwater scene.
@@
• Compared to state-of-the-art methods, our UVZ demonstrates more attractive results in both qualitative and quantitative comparisons. Moreover, UVZ exhibits generalization abilities to other visual tasks without parameter tuning.
II Related work
II-A Traditional Methods
@@
In response to the degradation characteristics of images, model-free methods notably enhance the contrast and saturation of the images. Ancuti et al. [15] fused color compensated and white balance images to obtain natural underwater image. Marques et al. [5] used contrast to guide the generation of images with salient details and brightness enhancement. Based on the concept of contour bougie morphology, Yuan et al. [6] enhanced the image quality by mitigating scattering blur and contrast stretching. Zhang et al. [3] introduced the attenuation characteristics to correct the color, and introduced the integral and squared integral maps to enhance the contrast. Kang et al. [21] decompose the image into mean intensity, contrast, and structure, the three components are fused in different ways to obtain the enhancement result. However, such methods overlook the characteristics of underwater degradation process, especially the depth between the object and the camera [22, 23], leading to color distortion or improper enhancement.
@@
Model-based methods explore the process of underwater image degradation and use priors and assumptions to model this process. Berman et al. [24] used the haze-lines model to enhance the attenuated colors. Alenezi et al. [17] introduced gradient and absorption differences associated with color channels into background light and transmission map estimation. Yuan et al. [7] enhanced the visibility of images by dark channel prior and color compensation in two color spaces, respectively. Zhou et al. [23] introduced light attenuation features and feature prior in estimating background light and transmission map. Chandrasekar et al. [25] combine underwater image formation models and depth information to remove backscatter, and utilize lightweight networks and retinex model to further enhance the image. Due to the variability of underwater environments, the parameters and priors set by these methods are limited and cannot fully adapt to specific underwater scenarios.
@@
II-B Deep Learning Methods
In recent years, deep learning methods have been widely applied in UIE due to remarkable effect. Wang et al. [26] introduced HSV space for the first time to adjust the brightness and saturation of underwater images. Jiang et al. [27] incorporated the laplacian pyramid to construct a lightweight multi-scale network, and improved the performance by recursive strategy and residual learning. Zhou et al. [18] combined RGB and HSV color spaces to enhance the underwater images, and used attention and color-guided map to guide the network to perceive critical information. Zhang et al. [28] introduced multi-level sub-networks to enrich the feature representation, and proposed a triple attention module to selectively cascade different features. By observing from the perspective of color spectrum distribution, Rao et al. [19] proposed a two-stage network using texture cues and color cues to enhance the underwater images. Wang et al. [29] integrated the wavelet transform and color compensation mechanism in the network to perform multi-level enhancement of details and colors.
To avoid the need for paired underwater images, unsupervised learning methods have received considerable attention. Jiang et al. [30] migrated image dehazing from air to the underwater domain, thus avoiding the reliance of paired underwater images. Liu et al. [11] constructed a twin adversarial contrastive learning framework and utilized underwater object detection tasks to guide image enhancement. Jiang et al. [31] employed global-local adversarial mechanism to supervise image realism. Li et al. [32] introduced comparative learning into UIE and utilized the discriminator to evaluate the quality of different regions.
II-C Depth-related Methods
The widespread application of depth information in image dehazing [33, 34, 35] has enlightened the understanding of underwater scenes in the UIE field. Zhou et al. [36] used channel intensity prior and unsupervised method to generate two depth maps, further eliminated the blurring problem caused by scattering effects. Chen et al. [20] introduced position attention and dilated convolution into the depth estimation network, guiding weight allocation by the predicted depth map. Fu et al. [12] reconstructed the raw image by predicting a depth-related transmission map, thereby obtaining the scene radiance for degradation. Cong et al. [13] used the depth map and the physical model for image restoration, with dual-discriminators constraining the generation of the restored image and the depth map. Hou et al. [37] improved the estimation of transmission map by the local background light and reflectance map, which can eliminate the influence of non-uniform light in underwater scenes.
III Methodology
As shown in Fig. 2, UVZ constitutes a two-stage UIE framework. First, we input the raw image $X\in\mathbb{R}^{H\times W\times 3}$ to DEN with a standard encoder-decoder architecture. Within this process, a dual-attention module (DAM) is introduced to focus on severely degraded regions, thereby obtaining a depth map $d\in\mathbb{R}^{H\times W\times 1}$ with well-defined near-far regions. During the training process, we generate a regression map $\widehat{X}\in\mathbb{R}^{H\times W\times 3}$ for color restoration via ASN to promote greater correlation between $d$ and $X$. To refine the restoration of $\widehat{X}$, ASN fuses the decoded features from DEN with residual skip blocks (RSB). Second, we design a depth perception module (DPM) for near-far region learning in DGEN, consisting of two branches: swin transformer [38] and multi-kernel convolution. The two branches are used for non-local and local feature learning, respectively, with selective fusion guided by $d$. In response to DPM, we apply an R3N transformation on $d$, as presented in Fig. 3. Using the adjusted depth maps and the learnable parameters, we constructed a fusion strategy with receptive field adaptation for a more reasoned combination of non-local and local features.
III-A Depth Perception Module
For most UIE networks, convolution is commonly used to capture local information over short distances. While these methods effectively eliminate most color deviation, certain regions may still be affected by blue-green, particularly in far regions with severe degradation. As the enhancement process of far point is mainly dominated by the surrounding degraded features, local feature extraction instead hinders the restoration of underwater colors. To locate regions with severe degradation, we notice the correlation between depth and degradation in the underwater imaging model. A reasonable depth map can assess the quality of different regions to a certain extent, thereby guiding a more comprehensive color enhancement. To this end, we constructed a DPM that guides perception with depth, and adaptively enhance different regions through non-local and local learning branches. DPM is illustrated in Fig. 2.
@@ III-A Depth Perception Module
Figure 3: R3N transformation of a depth map. The input is the depth map d and the outputs are the feature maps $d_{rev}$, $d_{1}$, $d_{3}$, and $d_{5}$. The process within the dashed line is a 3 × 3 region normalization (RN).
@@
To adapt the receptive fields of the two branches, we design an R3N transformation for the depth map $d$, as presented in Fig. 3. First, we reshape $d$ to $d_{1}\in\mathbb{R}^{h\times w\times 1}$ consistent with the encoded feature scale, which is performed by nearest interpolation. Second, we alter the weight distribution of the near and far scenes by reversing the depth map $d_{rev}=1-d$, so that the non-local branch focuses more on high-quality information. In addition, through the idea of region normalization (RN) [39], we smooth the weight distribution of the degraded regions, thereby avoiding the interference of far points with higher degradation on local information capture. RN divides the regions based on the size of the convolution kernel in the local branch, calculating the mean value of each region for normalization. The specific process is as follows:
@@
where $R_{1},R_{2},\ldots,R_{n}$ represent the different division regions, $n=(h\texttimes w)/k^{2}$. $x_{ij}$ represents the jth pixel of region $R_{i}$, and all pixels in that region are represented by the mean value $\mu_{i}$. Ultimately, we employ RN to obtain $d_{3}$ and $d_{5}$ corresponding to 3 × 3 and 5 × 5 convolutions. By guiding the selective fusion in DPM, the transformed depth map will enable the network to develop a reasonable perception of the degraded regions, making the enhanced colors more natural.
We adopt the swin transformer to implement non-local branch for capturing dependencies. Compared with the traditional transformer structure, the swin transformer utilizes the shifted windows to fully enhance the receptive field, further extending the non-local interaction ability. The input encoded features are first partitioned into a non-overlapping patch sequence $f_{i}=[p_{1},p_{2},\ldots,p_{m}]$, then flattened into linear vectors and computed as follows:
@@
On the other hand, local information should play a crucial role in near regions with good imaging. Compared to the non-local region, adjacent features tend to be more helpful for the enhancement process in that region. In the local branch, we propose a local feature extractor based on multi-kernel convolutions, including four 3 × 3 convolutions, four residual blocks, and three convolutions with different kernel sizes (1, 3, and 5), where the activation functions are LeakyReLU. To improve the spatial distribution perception of underwater scenes, multi-kernel convolutions are applied to obtain more perceptual fields, thereby enhancing scene details of varying sizes in multi-scale encoding features. We multiply the convolutional features $f_{c1}$,$f_{c3}$,$f_{c5}$ with the corresponding depth maps $d_{1}$, $d_{3}$, $d_{5}$, and dynamically adjust the weights by three learnable parameters $\beta_{1}$, $\beta_{3}$, and $\beta_{5}$:
@@
III-B Dual-Attention Module

### v3

@@ Abstract
The underwater images often suffer from color deviations and blurred details. To address these issues, many methods employ networks with an encoder/decoder structure to enhance the images. However, the direct skip connection overlooks the differences between pre- and post-features, and deep network learning introduces information loss. This paper presents an underwater image enhancement network that focuses on pre-post differences. The network utilizes a multi-scale input and output framework to facilitate the underwater image enhancement process. A novel cross-wise transformer module (CTM) is introduced to guide the interactive learning of features from different periods, thereby enhancing the emphasis on detail-degraded regions. To compensate for the information loss within the deep network, a feature supplement module (FSM) is devised for each learning stage. FSM merges the multi-scale input features, effectively enhancing the visibility of underwater images. Experimental results across several datasets demonstrate that the integrated modules yield significant enhancements in network performance. The proposed network exhibits outstanding performance in both visual comparisons and quantitative metrics. Furthermore, the network also exhibits good adaptability in additional visual tasks without the need for parameter tuning. The code and results are released in https://github.com/WindySprint/UVZ.
keywords:
Underwater Image Enhancement, Depth-guided Perception, Near-far Scenarios, Color Correction
journal: Nuclear Physics B
[…]
1 Introduction
@@
In the early days of UIE field, researchers have proposed model-free and model-based traditional methods. Model-free methods [5, 15, 16] enhances underwater images by modifying pixels in degraded channels and blurred regions. On the other hand, model-based methods [7, 17] introduce physical cues to build the underwater degradation process, and reverse it to restore underwater images. In recent years, deep learning methods [11, 18, 19] have shown remarkable enhancements, leveraging benchmark data and miscellaneous networks. Despite significant progress achieved with the above methods, challenges emerge when addressing widely varying degrees of regional degradation. As shown in Fig. 1, most UIE methods have considerable visual improvement in the near scene, but still retain notable color deviation in the far scene. Fig. 2 highlights the significant differences between near and far scenes, thereby underscoring the motivation behind this work. A more detailed description is provided in Section 3.1.
Figure 2: Motivation of our proposed depth-guided perception. The degradation degree varies significantly between the near scene and far scene. In comparison, the visual quality of the former and the adjacent regions exhibits relatively better performance. For the two scenes, we should implement different feature learning strategies.
For more precise color enhancement, we believe that the integration of depth information related to scene distance is essential, which benefits the region division in underwater images. Previous works [12, 13, 20] attempted to incorporate depth information related to scene distance, contributing to the response of degraded regions. However, direct weight adjustments may restrict the exploration of visual cues, as the adjacent features of far and near points exhibit varying gains in the enhancement process. In near regions of relatively good quality, local information from surrounding features is critical to the quality of the enhanced image, whereas in the far regions of heavily degradation, non-local information from other regions can provide more valuable insights. With the above analysis, we aim to construct an adaptive region enhancement method in underwater scenes, achieving comprehensive visual enhancement by differentiating in utilizing local and non-local information. To achieve this goal, there are two main challenges:(1) On the one hand, the local feature capture capability helps to enhance the near region where the surrounding information is enough. On the other hand, the remote dependency establishment capability helps to enhance the far region where the surrounding information is lost. However, due to feature space differences, there is mutual interference between these two capabilities. (2) Diverse underwater scenes tremendously raise the difficulty of predicting the depth map.
@@
Specifically, we propose a novel UIE method based on depth-guided perception, UVZ, which includes depth estimation network (DEN), auxiliary supervision network (ASN), and depth-guided enhancement network (DGEN). Firstly, the proposed model generates a depth map by DEN, and uses ASN to further supervise the rationality of the depth map during training. When enhancing each point, the depth information is used to determine the contribution of non-local and local information. Meanwhile, the desired features for different regions are captured by non-local and local branches in DGEN. To adapt to varying receptive fields, we design an R3S transformation for the estimated depth map, guiding feature fusion in different regions. Through the above process, UVZ effectively and comprehensively enhances the color, contrast, and clarity of underwater images.
@@
• We design a novel depth-guided perception framework, UVZ. For non-homogeneous degraded underwater scenes, UVZ employs the adapted depth maps to implement a learning strategy that adaptively incorporates non-local and local features.
@@
• To enhance the ability to distinguish degraded regions, we incorporate an attention estimation-scene reconstruction training process, aiming to strengthen the connection between depth maps and underwater scenes.
@@
• Compared to state-of-the-art methods, UVZ demonstrates more attractive results across extensive qualitative and quantitative comparisons. Moreover, UVZ exhibits generalization abilities to other visual tasks without parameter tuning.
2 Related work
2.1 Traditional Methods
@@
In response to the degradation characteristics of images, model-free methods notably enhance the contrast and saturation of the images. The early studies focused on adjusting image pixel values in the spatial domain. Ancuti et al. [15] fused color compensated and white balance images to obtain natural underwater image. Marques et al. [5] used contrast to guide the generation of images with salient details and brightness enhancement. As the research progressed, some methods introduced morphology and structural properties to understand underwater images. Based on the concept of contour bougie morphology, Yuan et al. [6] enhanced the image quality by mitigating scattering blur and contrast stretching. Kang et al. [21] decompose the image into mean intensity, contrast, and structure, the three components are fused in different ways to obtain the enhancement result. Additionally, Zhang et al. [3] introduced the attenuation characteristics to correct the color, and introduced the integral and squared integral maps to enhance the contrast. Overall, model-free methods overlook the characteristics of underwater degradation process, especially the depth between the object and the camera [22, 23], leading to color distortion or improper enhancement.
@@
Model-based methods explore the process of underwater image degradation and use priors and assumptions to model this process. Alenezi et al. [17] introduced gradient and absorption differences associated with color channels into background light and transmission map estimation. Zhou et al. [23] introduced light attenuation features and feature prior in estimating background light and transmission map. Chandrasekar et al. [24] combine underwater image formation models and depth information to remove backscatter, and utilize lightweight networks and retinex model to further enhance the image. In addition, some methods introduced prior knowledge from other domains to improve visual quality. Berman et al. [25] used the haze-lines model to enhance the attenuated colors. Yuan et al. [7] enhanced the visibility of images by dark channel prior and color compensation in two color spaces, respectively. Due to the variability of underwater environments, the parameters and priors set by such methods are limited and cannot fully adapt to specific scenarios.
2.2 Deep Learning Methods
In recent years, deep learning methods have made significant progress in UIE tasks. Classical methods were devoted to optimize the network structure to achieve better visual enhancement. Jiang et al. [26] incorporated the laplacian pyramid to construct a lightweight multi-scale network, and improved the performance by recursive strategy and residual learning. Zhang et al. [27] introduced multi-level sub-networks to enrich the feature representation, and proposed a triple attention module to selectively cascade different features. To enrich the input features, some studies have introduced other color spaces. Wang et al. [28] introduced HSV space for the first time to adjust the brightness and saturation of underwater images. Zhou et al. [18] combined RGB and HSV color spaces to enhance the underwater images, and used attention and color-guided map to guide the network to perceive critical information. Furthermore, the researchers have delved into the structural composition of images from diverse perspectives of data distribution. By observing from the perspective of color spectrum distribution, Rao et al. [19] proposed a two-stage network using texture cues and color cues to enhance the underwater images. Wang et al. [29] integrated the wavelet transform and color compensation mechanism in the network to perform multi-level enhancement of details and colors.
To avoid the need for paired underwater images, unsupervised learning methods have received considerable attention. Among them, generative adversarial network (GAN) has been a prevalent framework. Jiang et al. [30] employed global-local adversarial mechanism to supervise image realism. Li et al. [31] introduced comparative learning into UIE and utilized the discriminator to evaluate the quality of different regions. Beyond GAN, other learning methods such as transfer learning and comparative learning also show immense potential. Jiang et al. [32] migrated image dehazing from air to the underwater domain, thus avoiding the reliance of paired underwater images. Liu et al. [11] constructed a twin adversarial contrastive learning framework and utilized underwater object detection tasks to guide image enhancement.
2.3 Depth-related Methods
The widespread application of depth information in image dehazing [33, 34, 35] has enlightened the understanding of underwater scenes in the UIE field. Traditional methods incorporated depth estimation to mitigate underwater degradation effects on image quality. Zhou et al. [36] used channel intensity prior and unsupervised method to generate two depth maps, further eliminated the blurring problem caused by scattering effects. Hou et al. [37] improved the estimation of transmission map by the local background light and reflectance map, which can eliminate the influence of non-uniform light in underwater scenes. Simultaneously, deep learning methods utilized depth information to guide the scene weights for targeted distribution. Chen et al. [20] introduced position attention and dilated convolution into the depth estimation network, guiding weight allocation by the predicted depth map. Fu et al. [12] reconstructed the raw image by predicting a depth-related transmission map, thereby obtaining the scene radiance for degradation. Cong et al. [13] used the depth map and the physical model for image restoration, with dual-discriminators constraining the generation of the restored image and the depth map.
3 Methodology
3.1 Overall Architecture
Due to the regional degradation gap in underwater scenes, local dependence for perceiving surrounding features against remote dependence for perceiving global structures, it is nontrivial to learn the two perceptual capabilities from the diverse samples. The existing UIE methods eliminate most color deviation, but certain regions may still be affected by blue-green, particularly in far-filed regions with severe degradation. Because the enhancement process of far point is mainly dominated by the surrounding degraded features, and the local information instead hinders the underwater color correction. The remote dependence considers the overall scene from a macroscopic perspective, but lacks a precise definition and identification of the degree of regional degradation. To locate regions with severe degradation, we note the correlation between depth and degradation in underwater imaging models. A reasonable depth map can provide a intuitive assessment of the image quality in different regions, which is important for fusing local and global perceptual capabilities. In this work, we focus on building an integrated UIE framework guided by physical cues, aiming to achieve comprehensive color enhancement for diverse underwater degraded scenes.
As shown in Fig. 3, UVZ constitutes a two-stage UIE framework. In the first stage, we input the raw image $X\in\mathbb{R}^{H\times W\times 3}$ to the depth estimation network (DEN), and output the depth map $d\in\mathbb{R}^{H\times W\times 1}$ through encoding-decoding and two scale changes. Within the process of feature mapping, a dual-attention module (DAM) is introduced to monitor severely degraded regions from channel and spatial perspectives, allowing for a more explicit definition of the near and far regions. By introducing scene reconstruction and cyclic consistency loss, CycleGAN [38] ensures consistency between the target images and the source images. Inspired by this, we introduce the auxiliary supervision network (ASN) to generate the regression map $\widehat{X}\in\mathbb{R}^{H\times W\times 3}$ in the training phase, and jointly supervise the losses of DEN and ASN to promote greater correlation between $d$ and $X$. To refine the restoration of $\widehat{X}$, ASN fuses the decoded features from DEN via residual skip blocks (RSB), which further promotes the rationality of $\widehat{X}$ on the overall scene.
@@ 3.1 Overall Architecture
In the second stage, we design a depth perception module (DPM) for near and far feature learning. DPM consists of two distinct branches for non-local and local feature learning, with selective fusion guided by $d$ from the previous stage. In response to the DPM, we design an R3S transformation for $d$, as shown in Fig. 4. Using the adjusted depth maps and the learnable parameters, the depth-guided enhancement network (DGEN) constructs a fusion strategy with receptive field adaptation to combine non-local and local features more rationally. Meanwhile, DGEN incorporates multiple residual blocks (RB) to solve the gradient vanishing. Both RB and RSB extract features through convolution-activation-convolution structures and residual connection, while the latter includes an additional 1 × 1 convolution block for fusion.
3.2 Depth Perception Module
@@
Figure 4: R3S transformation of a depth map. The input is the depth map d and the outputs are the feature maps $d_{rev}$, $d_{1}$, $d_{3}$, and $d_{5}$. The process within the dashed line is a 3 × 3 region smoothing (RS).
@@
We constructed a DPM that guides perception with depth, and adaptively enhance different regions through non-local and local learning branches. Specifically, to adapt the receptive fields of the two branches, the DPM applies an R3S transformation to the depth map $d$. As shown in Fig. 4, we first reshape d to $d$ to $d_{1}\in\mathbb{R}^{h\times w\times 1}$ consistent with the encoded feature scale, which is performed by nearest interpolation. Second, we alter the weight distribution of the far and near scenes by reversing the depth map $d_{rev}=1-d$, so that the non-local branch focuses more on high-quality information. In addition, we introduce a region smoothing (RS) process to remove noise from degraded regions, thereby avoiding the interference of far points with higher degradation on local information capture. RS divides the regions based on the size of the convolution kernel in the local branch, calculating the mean value of each region for normalization. The specific process is as follows:
@@
where $R_{1},R_{2},\ldots,R_{n}$ represent the different division regions, $n=(h\texttimes w)/k^{2}$. $x_{ij}$ represents the jth pixel of region $R_{i}$, and all pixels in that region are represented by the mean value $\mu_{i}$. Ultimately, we employ RS to obtain $d_{3}$ and $d_{5}$ corresponding to 3 × 3 and 5 × 5 convolutions. By guiding the selective fusion in DPM, the transformed depth map will enable the network to develop a reasonable perception of the degraded regions, making the enhanced colors more natural.
We adopt the Swin Transformer [39] to implement non-local branch for capturing dependencies. Compared to the Vision Transformer, Swin Transformer captures remote dependencies at diverse scales by constructing hierarchical feature mappings, which allows it to capture details of different shapes (e.g., fish, plants, and stones, etc.) in underwater scenes. Additionally, Swin Transformer introduces a shifted windows mechanism to enhance the receptive field, further extending the ability to facilitate non-local feature interactions. The input encoded features are first partitioned into a non-overlapping patch sequence $f_{i}=[p_{1},p_{2},\ldots,p_{m}]$, then flattened into linear vectors and computed as follows:
@@
On the other hand, local information should play a crucial role in near regions with good imaging. Compared to the non-local region, adjacent features tend to be more helpful for the enhancement process in that region. In the local branch, we propose a local feature extractor based on multi-kernel convolutions, including four 3 × 3 convolutions, four residual blocks, and three convolutions with different kernel sizes (1, 3, 5), where the activation functions are LeakyReLU. To improve the spatial distribution perception of underwater scenes, multi-kernel convolutions are applied to obtain more perceptual fields, thereby enhancing details of varying sizes in multi-scale encoding features. We multiply the convolutional features $f_{c1}$,$f_{c3}$,$f_{c5}$ with the corresponding depth maps $d_{1}$, $d_{3}$, $d_{5}$, and dynamically adjust the weights by three learnable parameters $\beta_{1}$, $\beta_{3}$, and $\beta_{5}$:
@@
3.3 Dual-Attention Module